# Radif Corpus: A Symbolic Dataset for Non-metric Iranian Classical Music

### v1

@@ Abstract
In this study, we introduce the first digital corpus representing the complete non-metrical radif repertoire, covering all 13 existing components of this repertoire. We provide MIDI files (about 281 minutes in total) and data spreadsheets describing notes, note durations, intervals, and hierarchical structures for 228 pieces of music. We faithfully represent the tonality including quarter-tones, and the non-metric aspect. Furthermore, we provide supporting basic statistics, and measures of complexity and similarity over the corpus.
@@ 1 Introduction
In this work, we take a step toward addressing this gap by introducing a dataset specifically focused on Iranian non-metric classical music, aiming to support and inspire future studies in this area. We begin by introducing Iranian classical music and its core repertoire, the radif. After reviewing previously published datasets, we present our own dataset in detail. Finally, we provide a statistical and visual overview of the dataset, which can serve as a useful reference for researchers and practitioners.
@@ 1.2 Exploring Datasets
To the best of our knowledge, there was no symbolic corpus available for Iranian music before our previous work, in which we introduced the Shour Corpus [20]. This corpus includes one section of the radif (shur) and was used in our study on discovering patterns and producing meaningful variations through grammatical representation (compression) in this musical style.
@@ 1.3 Our Contribution
At the time of writing this paper, to the best of our knowledge, there is no symbolic dataset covering the entire radif. This led us to create the Radif Corpus, which includes all non-metric pieces from Mīrzā ’Abdollāh’s radif. Out of the several transcriptions of Mīrzā ’Abdollāh’s radif, we have selected the edition titled “Radif Analysis - based on the notation of Mīrzā ’Abdollāh’s radif with annotated visual description” by Dariush Talai [25]. This edition consists of a recorded performance, together with a score derived from the performance, notated with hierarchical structure. Although radif also includes some metric pieces, which are usually performed at the end of each dastgāh/āvāz, our corpus excludes them as we are focusing on non-metric music. Figure 1 presents an example of a transcription from the book.
@@ 2 Radif Corpus Description
[…]

### v3

@@ Abstract
In this study, we introduce a digital corpus representing the complete non-metrical radif repertoire, covering all 13 existing components of this repertoire. We provide MIDI files (about 281 minutes in total) and data spreadsheets describing notes, note durations, intervals, and hierarchical structures for 228 pieces of music. We faithfully represent the tonality including quarter-tones, and the non-metric aspect. Furthermore, we provide supporting basic statistics, and measures of complexity and similarity over the corpus.
@@ 1 Introduction
In this work, we take a step toward by introducing a dataset specifically focused on Iranian non-metric classical music, aiming to support and inspire future studies in this area. We begin by introducing Iranian classical music and its core repertoire, the radif. After reviewing previously published datasets, we present our own dataset in detail. Finally, we provide a statistical and visual overview of the dataset, which can serve as a useful reference for researchers and practitioners.
@@ 1.2 Exploring Datasets
In [20], the authors introduce a MIDI dataset for the Mīrzā ’Abdollāh radif, which, to the best of our knowledge, is the first study of its kind in this field. Furthermore, in our previous work, we introduced the Shour Corpus [21], which is used in our study on discovering patterns and producing meaningful variations through grammatical representation (compression) in this musical style.
@@ 1.3 Our Contribution
A symbolic dataset encompassing the full radif along with multiple representations, hierarchical structure, and statistical metadata has been lacking in the literature. This led us to create the Radif Corpus, which includes all non-metric pieces from Mīrzā ’Abdollāh’s radif. Out of the several transcriptions of Mīrzā ’Abdollāh’s radif, we have selected the edition titled “Radif Analysis - based on the notation of Mīrzā ’Abdollāh’s radif with annotated visual description” by Dariush Talai [26]. This edition consists of a recorded performance, together with a score derived from the performance, notated with hierarchical structure. Although radif also includes some metric pieces, which are usually performed at the end of each dastgāh/āvāz, our corpus excludes them as we are focusing on non-metric music. Figure 1 presents an example of a transcription from the book.
@@ 2 Radif Corpus Description
Octaves. We consider the lowest note in the first gūsheh of each dastgāh/āvāz as the start note of the main octave. For example in the first gūsheh of our corpus the lowest note is F3 so the main octave is from F3 to F4. The music in the main octave is represented only by its symbols and accidental signs (if any are present). For notes in octaves other than the main one, we use ‘+’ or ‘-’ followed by a number to indicate the number of octaves above or below the main octave. For example, an Akk from one octave higher than the main octave would be written as Akk+1.